I need someone to teach me how this question work because I can't understand it.
In a single deal of 5 cards from a standard 52-card deck, what is the probability of being dealt 5 clubs?
I don't understand it at all.
Thank very much.

Found 2 solutions by Alan3354, edjones:
Answer by Alan3354(69443)

I need someone to teach me how this question work because I can't understand it.
In a single deal of 5 cards from a standard 52-card deck, what is the probability of being dealt 5 clubs?
-------------
The 1st card has a 13/52 chance (13 clubs out of 52 cards)
The 2nd is 12/51
3rd is 11/50
4th is 10/49
5th is 9/48
Overall, it's the product
= 13*12*11*10*9/52*51*50*49*48
= 33/66640
=~ 0.0004952

Answer by edjones(8007)

There are 13 clubs in 52 cards.
nCr=n!/((n-r)!r!) Combination of n things taken r at a time.
13C5/52C5 Combinations of 5 out of 13 cards / combination of 5 out of 52 cards
=1287/259860
=.000495... probability
.
Ed